OpenStudy (anonymous):

Write 3log6(x) + log6(y-2) -2log6(a) as a single log expression

OpenStudy (anonymous):

\[\log6^{3}x ^{3}+\log(6y-12)-\log6^{2}a ^{2}\] \[\log \frac{ 216x ^{3}\times(6y-12) }{36a ^{2} }\] \[\log \frac{ 1296x ^{3}y-2592 }{36a ^{2}}\] \[\log \frac{ 36x ^{3}y-72 }{a ^{2}}\] I think is this.=)

OpenStudy (anonymous):

where did you get this from?

OpenStudy (anonymous):

i do it just now

OpenStudy (anonymous):

but you got it from a website

OpenStudy (anonymous):

u know the logarithm theorem? Use the theorem. From website?

OpenStudy (noelgreco):

I think the 6 is the base of the log.

OpenStudy (anonymous):

You can make it simpler: \[\begin{align} &3\log_6 x + \log_6 (y-2)-2\log_6 a\\ &\log_6\left(\frac{x^3(y-2)}{a^2}\right) \end{align}\]

OpenStudy (anonymous):

oh~i mistaken~i think the 6 is multiple~